OpenStudy (anonymous):

tell wether the expression is equivalent (12-7) + x and 5x

OpenStudy (anonymous):

12-7=5

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

So what do u think 5+x and 5x are qual or not?

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

What if x=5/4

OpenStudy (anonymous):

i have no clue ?

OpenStudy (anonymous):

im not good at math

OpenStudy (anonymous):

Ok..... if x=5/4 then